# Leveraging AI-Generated Emotional Self-Voice to Nudge People towards their Ideal Selves

### v1

@@ 3. Generated Emotional Self Voice for Nudging towards Ideal Self
In the context of HCI, it is useful to distinguish between self-voice and other-voices, such as those of other humans or robotic voices. In this work, we use the phrase “self-voice” to refer to authentic and generated versions of one’s own voice, and “Emotional Self Voice” (ESV) to refer to our proposed system and its generated outputs. We designed Emotional Self Voice to allow individuals to hear an ideal version of themselves responding to a situation in which they had either failed to achieve a personal goal, or were struggling to persist in achieving some goal. These responses aim to nudge individuals toward their ideal selves by presenting a vivid mental picture of what that self might sound like. In the following subsections, we describe the implementation of the Emotional Self Voice system in further detail.
@@
3.1. System Implementation
3.1.1. Step 1: Text Generation of Ideal Self Responses
We used an LLM to generate ideal-self responses based on a scenario provided by the user. The goal of the text generation step is two-fold: (1) to represent the personality and values of one’s ideal self, and (2) to generate cues for emotionally expressive speech. We first asked participants to describe scenarios in which they had failed to achieve a personal goal, or were struggling to establish a new personal habit. Then, we asked them about the characteristics of their ideal self in the provided scenario. In describing the scenario and characteristics of their ideal self, users essentially completed an exercise akin to a journaling prompt for self-reflection. However, the LLM-generated text exposes the user to possible responses that they might have not thought of. Finally, we also prompted the LLM to add vocal bursts (such as sigh), discourse markers (such as ”well”, ”I mean”, ”uh”), and other natural emotive embellishments to the generated responses to facilitate the generation of natural and emotionally expressive speech in the next step of our system.
The full prompt can be found in the Appendix.
@@ 3.1.2. Step 2: Emotional Speech Generation from Text
The emotional content of speech comes not only from its linguistic content but also how it is spoken — from pitch and intonations to non-verbal expressions (Cutler et al., 1997). Hence, in this step, we use Hume AI’s empathic voice interface [^3] https://dev.hume.ai/docs/empathic-voice-interface-evi/overview to generate emotionally expressive speech based on the output of Step 1. Although Hume AI and new text-to-speech models (TTS) (Wu et al., 2024) represent an advancement over previous TTS models that were only able to generate monotonic or emotionally neutral outputs [^4] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/, at the time of writing, the model’s API does not allow one to directly adjust the emotional parameters of the speech output. Thus, the emotional quality of the generated speech largely depends on the expressiveness of the text. In a later section, we present the evaluation of the emotional quality of the generated speech, which shows the reliability and expressiveness of the generation.
@@ 3.1.3. Step 3: Voice Cloning and Generating Emotional Self-Voice with Speech-to-Speech models
Increasingly, there are many open-source models (Kang et al., 2022; Betker, 2023; Wang et al., 2023) and commercial models (COQUE, 2023; Labs, 2024) that allow zero-shot and speech-to-speech generation with a target voice. In choosing the model to perform the voice cloning and speech-to-speech generation, we prioritize stability and speed of the generation, since the focus of this work is on the effect of an ESV system on individuals perception and behavior. Most open source models are unfortunately both computationally expensive and slow while the output quality is not convincing. In the end, we chose ElevenLab’s voice cloning API given its speed and quality.
We experimented with different text for the users to read for voice cloning, and we referenced best practices to achieve the best generation quality [^5] https://elevenlabs.io/docs/voices/voice-lab/instant-voice-cloning. Our initial tests with neutral text produced a more monotone voice, therefore, we settled on a set of emotionally positive and negative texts to elicit a broad range of emotional expression (Appendix).
Once a reference of the user’s voice is obtained, the next step is to create a transformation between the source speech (i.e., generated emotional speech from previous component) and the target voice (i.e., the user’s voice). Initially, we also experimented with ElevenLab’s text-to-speech API [^6] https://elevenlabs.io/docs/api-reference/text-to-speech and used the generated text from Step 1. The result yielded monotonic speech and the quality varied depending on the emotional tone of the target user voice. Thus, we use the emotionally generated speech from Step 2 as the source speech input to speech-to-speech API. We also tuned different parameters of the API such that the output preserves both the emotional range of the source speech and the characteristics of the target voice.
@@ 3.1.4. Final System Setup
Putting everything together, the final system implementation is as follows: we used ElevenLabs’ instant voice cloning API [^7] https://elevenlabs.io/docs/voices/voice-lab/instant-voice-cloning to synthesize a clone of self-voice clips based on short voice samples of about 1 minute, which returns a unique voice ID. We then used GPT-4o [^8] https://platform.openai.com/docs/models/gpt-4o (temperature=0.15) to generate a text response based on a given scenario and set of personalities. In order to generate emotionally expressive speech output, we use Hume AI’s LLM (EVI v1) to perform text-to-speech [^9] https://dev.hume.ai/docs/empathic-voice-interface-evi/overview. We set the gender of the voice configuration to match the gender of the user. The generated voice clip is then inputted as the source speech to ElevenLab’s speech-to-speech model (Multilingual v2, stability=1.0, similarity_boost=0.7, style=0) [^10] https://elevenlabs.io/docs/speech-synthesis/models to translate the generated speech into the user’s voice (target voice). The end-to-end generation time is about 30 seconds/sentence from generating the input text to generating the emotional self voice.
@@ 4.1. Scenarios
We chose scenarios within self-development, where individuals set out to achieve some goal. We were not considering situations where ESV could help interpersonal relationships. Although important, for this study, we chose not to focus on situations where other individuals’ perception of an event has a significant effect (e.g., asking the boss for a raise, where the relationship between the boss and the individual is unknown). We focused on situations where the individual’s perception of the event has a significant effect on consequential decisions and behaviors (e.g., having self-deprecating thoughts after a failure or a lack of courage to pursue a goal due to imposter syndrome). In particular, we looked at two types of scenarios within the category of individual goal-achievement: (1) overcoming failure to accomplish a goal, and (2) struggling to persist in having a personal habit, which are common challenges that occur in the process of achieving a goal. For each type of scenario, we asked participants to provide a personal scenario based on our prompts so that the scenarios elicit real emotional responses. We also asked participants to describe their ideal self characteristics in text. The scenarios and the characteristics were then provided to the LLM to generate a hypothetical ideal self response to the scenarios. Next, we describe the prompt we used for the human participants and the LLM for generating the text.
4.1.1. Scenario 1: Overcoming the Failure of Achieving a Goal
In this scenario, we prompted participants to reflect and write about a time when they had a goal or something they wanted to accomplish but they didn’t succeed. In particular, we asked: “What was the goal?”, “Why did you want to achieve it?”, “What was the reason you didn’t succeed?”, and “How did that make you feel?” In this situation of failure, one would reasonably feel discouraged, and worse, people with low self-compassion may attribute failure as a characteristic of themselves. The latter could exaggerate the severity of the event and deter future attempts at the goal. We adapted the common cognitive-behavior therapy strategy of reframing by replacing possible unhelpful thoughts with ones that bring out positive outlooks and constructive next steps (Wenzel, 2017). The prompt we used for the LLM to generate the ideal self text response is: “Your task is to imagine yourself as the person with these trait personalities would say to themselves in the given scenario that would help them reframe the situation and be resilient in the face of failure.”
4.1.2. Scenario 2: Struggling to Persist in Pursuing a Goal
We prompted participants to describe a personal habit that they are having a difficult time establishing. In particular, we asked: “What is the habit?”, “Why do you want to cultivate it?”, “What is the reason it is challenging?”, and “How do you think successfully forming the habit will make you feel?” In this scenario, the individual needs motivation, confidence, and commitment to persist. Research has shown that individuals who think of goals as identities felt the goals were easier to pursue and pursued goal-consistent choices (Dominick and Cole, 2020). We adapted the concept of ”goals as identities” where the ideal self embodies someone with the goal already integrated as their identity. The prompt we used for the LLM to generate the ideal self text response is: “Your task is to imagine yourself as the person with these trait personalities would say to themselves in the given scenario that would encourage them to keep up with the habit. You should try to embody the person when their habit has become their identity. Use the template of “I am a xxx person. ”
@@ 4.1.2. Scenario 2: Struggling to Persist in Pursuing a Goal
The motivation behind generating the text prompt is that people often might lack words to describe how they would ideally respond. Similar to how people confide in therapists or friends to hear how others might respond to their situation, providing generated responses might spark some inspiration.
For the full prompt to the LLM for both scenarios, please see Appendix.
4.2. Conditions
We are interested in how different aspects of the ESV system contribute to one’s perception of an event. We employed three conditions: Control, Text, and ESV. The Control condition is a passive control that is meant to mirror the internal reflection process one might have on their own or prompted by someone like a therapist or a friend. In this condition, the participants do not receive AI-generated text responses based on their ideal self and scenario, and they do not hear the generated self-voice saying the responses. The Text condition is an active control condition that isolates the effect of self-voice. In this condition, the participants receive generated text responses but do not hear the generated self-voice saying the responses to them. Finally, the ESV condition goes through the voice cloning procedure and receives the generated response in both text and self-voice. The study was designed to be between-condition to avoid the spill-over effect of experiencing multiple modalities of the intervention.
4.3. Participants
We recruited 64 participants from Prolific, and 4 were dropped from the data pool due to failing attention checks and giving incomplete responses. This leaves us with a final set of 60 participants equally split across three conditions. The participants group consists of 28 male, 30 female, and 2 non-binary individuals, with a mean age of 35.72 (SD = 8.93). All participants resided in the United States. They were screened for being at least 18 years old and speaking fluent English. Participants in the ESV condition were screened for having normal or corrected to normal hearing. The participants spent an average of 32.4 minutes on the study, and the participants received \$15 for completing the study. The study was approved by our Institutional Review Board (protocol #2408001387).
Figure 4. An overview of the user study protocol.
@@ 4.4.1. Onboarding
Participants provided informed consent to voice cloning and to have their survey data used anonymously prior to proceeding to the rest of the survey. Once enrolled in the study, participants provided their gender and age, and completed a questionnaire measuring Alexithymia which assesses difficulty in perceiving emotions (PAQ-S) (Preece et al., 2023), the Plymouth Sensory Imagery Questionnaire which assesses the vividness of mental image (Psi-Q sound and feeling subscales (Andrade et al., 2014)), and the Self-Compassion Scale (SSCS-S (Neff et al., 2021)). They also completed the Mini-IPIP (Donnellan et al., 2006) (assesses the Big Five personality dimensions) twice, with respect to who they actually are and who they would ideally like to be. We then used the contrast of the IPIP scores of the ideal and current self to calculate the amount of self-discrepancy (Stanley and Burrow, 2015).
@@ 4.4.2. Voice Cloning
Next, participants in the ESV condition were asked to record an emotionally expressive sample of their voice by reading sentences aloud (see Appendix). They then listened to their true voice recording and their cloned self-voice reading an emotionally neutral text. They were asked about their perception of hearing their own voice and then rated the quality of the cloned voice in terms of familiarity, similarity, naturalness, and quality. The full questionnaire is in the Appendix.
@@ 4.4.3. Intervention
Participants were presented with a reflection task related to the two scenarios in a randomized order, counterbalanced within each condition. For each scenario, they were first asked to describe a scenario by answering questions described in Section 4.1. During the intervention, all participants were asked to “write down 5 adjectives that describe your ideal self in facing the scenario you described.” After inputting the ideal-self attributes, the control condition group was simply asked to “imagine this ideal version of yourself with these personalities and what would your ideal self say to yourself in this scenario.” Both Text and ESV conditions received generated text by the LLM, and the ESV condition group also received the cloned self-voice saying the generated text. In addition, participants in these two conditions also were presented with two sliders that allowed them to adjust the generated responses a maximum of three times. The two sliders were positivity and emotional expressiveness on a scale of -3 to +3. These two scales are common features of sentiment. The intention behind the sliders is two-fold: first, the participants might not be immediately satisfied with generated responses compared to how they envisioned what their ideal self might say. We wanted to provide an opportunity for fine-tuning the response. Second, we are also interested in sampling people’s preferences to develop an emotional characterization of an ideal self. We were inspired by the Gibbs Sampling with People (GSP) approach (Harrison et al., 2020), where in each iteration the participant uses a slider to manipulate a stimulus dimension to optimize a given criterion. At the end of the intervention, all participants were asked questions about the vividness, similarity, and preference of their ideal self. We adapted the self-continuity questionnaire (FSCQ (Sokol and Serper, 2020)), where we replaced the ”future self” with ”ideal self”. We believe this adaption is valid as the ideal self is a particular version of the future self.

### v2

@@ 3. Generated Emotional Self Voice for Nudging towards Ideal Self
In this work, we use the phrase “self-voice” to refer to authentic and generated versions of one’s own voice, and “Emotional Self Voice” (ESV) to refer to our proposed system and its generated outputs. We designed Emotional Self Voice to allow individuals to hear an ideal version of themselves responding to a situation in which they had either failed to achieve a personal goal, or were struggling to persist in achieving some goal. These responses aim to nudge individuals toward their ideal selves by presenting a vivid mental picture of what that self might sound like. In the following subsections, we describe the implementation of the Emotional Self Voice system in further detail.
3.1. System Implementation
3.1.1. Step 1: Text Generation of Ideal Self Responses
The goal of the text generation step is two-fold: (1) to represent the personality and values of one’s ideal self, and (2) to generate cues for emotionally expressive speech. We first asked participants to describe scenarios in which they had failed to achieve a personal goal, or were struggling to establish a new personal habit. Then, we asked them about the characteristics of their ideal self in the provided scenario. We used GPT-4o, an LLM, to generate ideal-self responses based on a scenario provided by the user. In describing the scenario and characteristics of their ideal self, users essentially completed an exercise akin to a journaling prompt for self-reflection. However, the LLM-generated text exposes the user to possible responses that they might have not thought of. Finally, we also prompted the LLM to add vocal bursts (such as sigh), discourse markers (such as ”well”, ”I mean”, ”uh”), and other natural emotive embellishments to the generated responses to facilitate the generation of natural and emotionally expressive speech in the next step of our system. The full prompt can be found in the Appendix A.1.
@@ 3.1.2. Step 2: Emotional Speech Generation from Text
The emotional content of a speech comes not only from its linguistic content but also how it is spoken — from pitch and intonations to non-verbal expressions (Cutler et al., 1997). Hence, in this step, we used Hume AI’s empathic voice interface [^3] https://dev.hume.ai/docs/empathic-voice-interface-evi/overview to generate emotionally expressive speech based on the output of Step 1. Although Hume AI and new text-to-speech models (TTS) (Wu et al., 2024) represent an advancement over previous TTS models that were only able to generate monotonic or emotionally neutral outputs [^4] https://openai.com/index/chatgpt-can-now-see-hear-and-speak/, at the time of writing, the model’s API does not allow one to directly adjust the emotional parameters of the speech output. Thus, the emotional quality of the generated speech largely depends on the expressiveness of the text. In a later section, we present the evaluation of the emotional quality of the generated speech, which shows the reliability and expressiveness of the generation.
@@ 3.1.3. Step 3: Voice Cloning and Generating Emotional Self-Voice with Speech-to-Speech models
Increasingly, there are many open-source models (Kang et al., 2022; Betker, 2023; Wang et al., 2023) and commercial models (COQUE, 2023; Labs, 2024) that allow zero-shot and speech-to-speech generation with a target voice. In choosing the model to perform the voice cloning and speech-to-speech generation, we prioritize stability and speed of the generation, since the focus of this work is on the effect of an ESV system on individuals’ perception and behavior. Most open source models are unfortunately both computationally expensive and slow while the output quality is not convincing. In the end, we chose ElevenLab’s voice cloning API [^5] https://elevenlabs.io/docs/voices/voice-lab/instant-voice-cloning given its speed and quality.
We experimented with different text for the users to read for voice cloning, and we referenced best practices to achieve the best generation quality [^6] https://elevenlabs.io/docs/voices/voice-lab/instant-voice-cloning. Our initial tests with neutral text produced a more monotone voice, therefore, we settled on a set of emotionally positive and negative texts to elicit a broad range of emotional expressions (see Appendix A.2).
Once a reference of the user’s voice is obtained, the next step is to create a transformation between the source speech (i.e., generated emotional speech from the previous component) and the target voice (i.e., the user’s voice). Initially, we also experimented with ElevenLab’s text-to-speech API [^7] https://elevenlabs.io/docs/api-reference/text-to-speech and used the generated text from Step 1. The result yielded monotonic speech and the quality varied depending on the emotional tone of the target user voice. Thus, we use the emotionally generated speech from Step 2 as the source speech input for the speech-to-speech API. We also tuned different parameters of the API such that the output preserves both the emotional range of the source speech and the characteristics of the target voice.
@@ 3.1.4. Final System Setup
Putting everything together, the final system implementation is as follows: we used ElevenLabs’ instant voice cloning API [^8] https://elevenlabs.io/docs/voices/voice-lab/instant-voice-cloning to synthesize a clone of self-voice clips based on short voice samples of about 1 minute, which returns a unique voice ID. We then used GPT-4o [^9] https://platform.openai.com/docs/models/gpt-4o (temperature=0.15) to generate a text response based on a given scenario and set of personalities. In order to generate emotionally expressive speech output, we use Hume AI’s LLM (EVI v1101010https://dev.hume.ai/docs/empathic-voice-interface-evi/overview) to perform text-to-speech. We set the gender of the voice configuration to match the gender of the user. The generated voice clip is then inputted as the source speech to ElevenLab’s speech-to-speech model [^11] https://elevenlabs.io/docs/speech-synthesis/models (Multilingual v2, stability=1.0, similarity_boost=0.7, style=0) to translate the generated speech into the user’s voice (target voice). The end-to-end generation time is about 30 seconds/sentence from generating the input text to generating the emotional self voice.
@@ 4.1. Scenarios
We chose scenarios within self-development, where individuals set out to achieve some goal. We did not consider situations where ESV could help interpersonal relationships. Although important, for this study, we chose not to focus on situations where other individuals’ perception of an event has a significant effect, e.g., asking the boss for a raise, where the relationship between the boss and the individual is unknown. We focused on situations where the individual’s perception of the event has a significant effect on consequential decisions and behaviors (e.g., having self-deprecating thoughts after a failure, or a lack of courage to pursue a goal due to imposter syndrome). In particular, we looked at two types of scenarios within the category of individual goal-achievement: (1) overcoming a past failure of accomplishing a goal, and (2) struggling to persist towards establishing a personal habit, which are common challenges that occur in the process of achieving a goal.
While both scenarios involve challenges in behavior change, they differ fundamentally in terms of the following key aspects:
Temporal and cognitive focus—“Facing failure in achieving a goal” is inherently retrospective. It involves reflection on past actions and outcomes, often accompanied by emotions like regret or disappointment. The cognitive process focuses on resilience, evaluating what went wrong, and learning from past experiences. “Experiencing difficulty in establishing a habit” is forward-looking. It emphasizes ongoing effort and the anticipation of future behavior, with a cognitive focus on persistence and motivation to form routines (Wood and Neal, 2007).
Motivation—Goal failure triggers a corrective motivation, prompting individuals to re-evaluate strategies or even adjust their goals (Yeager and Dweck, 2020). Habit formation struggles elicit a maintenance motivation, focusing on repetition and consistency.
Outcome vs. process orientation—Goals are typically tied to specific outcomes, while habits are associated with the process itself and focus on automaticity and repetition. A failed goal implies an unmet outcome, whereas a struggling habit suggests an ongoing process that may still succeed with continued effort (Clear, 2018).
@@
For each type of scenario, we asked participants to provide a personal scenario based on our prompts so that the scenarios elicit real emotional responses. We also asked participants to describe their ideal self characteristics in the text. The scenarios and the characteristics were then provided to the LLM to generate a hypothetical ideal-self response to the scenarios. The motivation behind generating the text prompt is that people often might lack words to describe how they would ideally respond. Similar to how people confide in therapists or friends to hear how others might respond to their situation, providing generated responses might spark some inspiration. Figure 4 shows examples of scenarios, ideal self attributes provided by study participants and the corresponding ideal self response generated by our system.
We had an iterative approach to the prompts for generating the ideal-self response. There are two aspects we focused on. First, the perspective of the response is inspired by the cognitive-behavioral therapy practice, which we detail in the following sections. Second is the emotional expressiveness of the response. The generation of the audio output is based on the text prompt, of which the sentiment reflects the sentiment of the text response. For example, adding discourse markers makes the responses sound more natural and engaging as described in section 3.1.1.
Next, we describe the prompts we used for the human participants and the LLM for generating the text.
Figure 5. An overview of the user study procedure. PsQ-i: Plymouth Sensory Imagery Questionnaire; SSC-S: Self-Compassion Scale; IPIP: assesses the Big Five personality dimensions; BRT: Benchmark Resilience Tool; FSCQ: Future-Self Continuity Questionnaire.
@@
4.1.1. Scenario 1: Overcoming the Failure of Achieving a Goal
In this scenario, we prompted participants to reflect and write about a time when they had a goal or something they wanted to accomplish but they didn’t succeed. In particular, we asked: “What was the goal?”, “Why did you want to achieve it?”, “What was the reason you didn’t succeed?”, and “How did that make you feel?” In this situation of failure, one would reasonably feel discouraged, and worse, people with low self-compassion may attribute failure as a characteristic of themselves. The latter could exaggerate the severity of the event and deter future attempts at the goal. We adapted the common cognitive-behavior therapy strategy of reframing by replacing possible unhelpful thoughts with ones that bring out positive outlooks and constructive next steps (Wenzel, 2017). The prompt we used for the LLM to generate the ideal-self text response is: “Your task is to imagine yourself as the person with these trait personalities would say to themselves in the given scenario that would help them reframe the situation and be resilient in the face of failure.”
4.1.2. Scenario 2: Struggling to Persist in Pursuing a Goal
We prompted participants to describe a personal habit that they are having a difficult time establishing. In particular, we asked: “What is the habit?”, “Why do you want to cultivate it?”, “What is the reason it is challenging?”, and “How do you think successfully forming the habit will make you feel?” In this scenario, the individual needs motivation, confidence, and commitment to persist. Research has shown that individuals who think of goals as identities felt the goals were easier to pursue and pursued goal-consistent choices (Dominick and Cole, 2020). We adapted the concept of “goals as identities” where the ideal self embodies someone with the goal already integrated as their identity. The prompt we used for the LLM to generate the ideal-self text response is: “Your task is to imagine yourself as the person with these trait personalities would say to themselves in the given scenario that would encourage them to keep up with the habit. You should try to embody the person when their habit has become their identity. Use the template of “I am a xxx person. ”
For the full prompt to the LLM for both scenarios, please see Appendix A.1.
4.2. Conditions
We are interested in how different aspects of the ESV system contribute to one’s perception of an event. We employed three conditions: Control, Text, and ESV. The Control condition is meant to mirror the internal imagination process one might have on their own or prompted by someone like a therapist or a friend. In this condition, the participants do not receive AI-generated text responses based on their ideal self and scenario, and they do not hear the generated self-voice saying the responses. The Text condition isolates the effect of self-voice. In this condition, the participants receive generated text responses but do not hear the generated self-voice saying the responses to them. Finally, the ESV condition goes through the voice cloning procedure and receives the generated response in both text and self-voice. The study was designed to be between-condition to avoid the spill-over effect of experiencing multiple modalities of the intervention.
4.3. Participants
We recruited 64 participants from Prolific, and 4 were dropped from the data pool due to failing attention checks and giving incomplete responses. This left us with a final set of 60 participants equally split across three conditions. Note that while our sample size aligns with prior work investigating voice-based interventions (Chan et al., 2021; Kim and Song, 2024), and we acknowledge the exploratory nature of this study. Given the novelty of our specific research questions, there were no directly comparable studies from which to derive expected effect sizes. This work therefore serves as an initial investigation to estimate effect sizes and inform sample size planning for future confirmatory studies. The participants group consisted of 28 male (M), 30 female (F), and 2 non-binary (NB) individuals (Control: 10M, 9F, 1NB; Text: 8M, 11F; 1NB; ESV: 10M, 10F). Participants ranged from age 18 and 56 (Control: mean= 33.8, SD=7.53; Text: mean=38.35, SD=8.60; ESV: mean=35, SD=10.27). We provide a full breakdown of participants’ demographics and other control variables per condition in Appendix Section 7. All participants resided in the United States. They were screened for being at least 18 years old and speaking fluent English. Participants in the ESV condition were screened for having normal or corrected to normal hearing. The participants spent an average of 32.4 minutes on the study, and the participants received \$15 for completing the study. The study was approved by our Institutional Review Board (protocol #2408001387).
@@ 4.4.1. Onboarding
Participants provided informed consent to voice cloning and to have their survey data used anonymously prior to proceeding to the rest of the survey. Once enrolled in the study, participants provided their gender and age, and completed a questionnaire measuring Alexithymia which assesses difficulty in perceiving emotions (PAQ-S) (Preece et al., 2023), the Plymouth Sensory Imagery Questionnaire which assesses the vividness of mental image (Psi-Q sound and feeling subscales (Andrade et al., 2014)), and the Self-Compassion Scale (SSCS-S (Neff et al., 2021)). They also completed the Mini-IPIP (Donnellan et al., 2006) (assesses the Big Five personality dimensions) twice, with respect to who they actually are and who they would ideally like to be. We then used the difference of the IPIP scores of the ideal and current self to calculate the amount of self-discrepancy (Stanley and Burrow, 2015).
@@ 4.4.2. Voice Cloning
Next, participants in the ESV condition were asked to record an emotionally expressive sample of their voice by reading sentences aloud (Appendix A.2). They then listened to their true voice recording and their cloned self-voice reading an emotionally neutral text. They were asked about their perception of hearing their own voice and then rated the quality of the cloned voice in terms of familiarity, similarity, naturalness, and quality. The full questionnaire is in the Appendix B.
@@ 4.4.3. Intervention
Participants were presented with a reflection task related to the two scenarios in a randomized order, counterbalanced within each condition. For each scenario, they were first asked to describe a scenario by answering questions described in Section 4.1. During the intervention, all participants were asked to “write down 5 adjectives that describe your ideal self in facing the scenario you described.” After inputting the ideal-self attributes, the control condition group was simply asked to “imagine this ideal version of yourself with these personalities and what would your ideal self say to yourself in this scenario.” Both Text and ESV conditions received generated text by the LLM, and the ESV condition group also received the cloned self-voice saying the generated text. In addition, participants in these two conditions also were presented with two sliders that allowed them to adjust the generated responses a maximum of three times. The two sliders were positivity and emotional expressiveness on a scale of -3 to +3 (Figure 3). These two scales are common features of sentiment. The intention behind the sliders is two-fold: first, the participants might not be immediately satisfied with generated responses compared to how they envisioned what their ideal self might say. We wanted to provide an opportunity for fine-tuning the response. Second, we are also interested in sampling people’s preferences to develop an emotional characterization of an ideal self. We were inspired by the Gibbs Sampling with People (GSP) approach (Harrison et al., 2020), where in each iteration the participant uses a slider to manipulate a stimulus dimension to optimize a given criterion. At the end of the intervention, all participants were asked questions about the vividness, similarity, and preference of their ideal self. We adapted the self-continuity questionnaire (FSCQ (Sokol and Serper, 2020)), where we replaced “future self” with “ideal self”. We believe this adaption is valid as the ideal self is a particular version of the future self. They were also asked to describe the perceived emotions from the generated text and the generated audio (ESV condition only).